OpenStudy (anonymous):

3^2y+1-28(3^y)+9=0 Find real values of y

OpenStudy (anonymous):

\[3^{2y+1}-28(3^y)+9=0~~?\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[3(3^{2y})-28(3^y)+9=0\] Let \(x=3^{y}\), then \(x^2=(3^y)^2=3^{2y}\). So what you really have is a quadratic equation: \[3x^2-28x+9=0\] Find the solutions for \(x\). But, keep in mind you want a solution for \(y\). If you happen to get something like \(x=2\), then a solution would be \(3^y=2\) or \(y=\log_32\).

OpenStudy (anonymous):

ok thank you very much.

OpenStudy (anonymous):

yw